# Towards Anytime Optical Flow Estimation with Event Cameras

### v1

@@ III Methodology
In this paper, we propose, for the first time, an Event Anytime Flow Estimation (EVA-Flow) framework, as shown in Fig. 2, which achieves high accuracy, high frame rate, and low latency with supervision solely dependent on low frame rate optical flow. EVA-Flow is designed to overcome the frame rate limitations of existing event-based optical flow datasets and produce temporally-dense optical flow estimation with accuracy boosts compared to contemporary methods. We first put forward a Unified Voxel Grid Representation (UVG) in Sec. III-A, which generates event representations with low latency. These UVG bins are sequentially fed into EVA-Flow, where high-frame-rate optical flow estimations are generated with low latency using a Spatiotemporal Motion Recurrent (SMR) module, which is detailed in Sec. III-B. Then, the details of the loss function and the supervision regime are described in Sec. III-C. Finally, we propose a new criterion to evaluate the reliability of intermediate high-frame-rate optical flow in Sec. III-D.
III-A Event Representation: Unified Voxel Grid
@@
To achieve anytime event-based optical flow estimation, we propose an Unified Voxel Grid (UVG) with high-frame-rate representations. The event representation is crucial for the event-based model [43, 44]. Voxel Grid [17] is a commonly used event representation that utilizes the temporal dimension of event data, and its effectiveness for optical flow estimation tasks has been verified in many studies [17, 11, 39]. The principle of the Voxel Grid is to discretize the spatially sparse and temporally-continuous information of events by averaging over time, and then utilize bilinear interpolation in both spatial and temporal dimensions to obtain a tensor representation of the event data in a voxel form. The tensor has a dimension of $B{\times}H{\times}W$, where $B$ represents the number of time steps, and its size determines the sampling accuracy of the event data in the temporal dimension. Each channel of the Voxel Grid can be viewed as an event representation at a specific time. Given a set of events $\left\{\left(x_{i},y_{i},t_{i},p_{i}\right)\right\}_{i\in[1,N]}$ and $B$ bins to discretize the time dimension, the definition of Voxel Grid is depicted as follows:
@@
Figure 3: The proposed Unified Voxel Grid (UVG) compared with Voxel Grid [17]. The lines of different colors indicate the temporal scope of events encompassed in their respective bins, while the vertical axis of each line represents the interpolated weight of the events at the point in time. Various line colors are employed to differentiate among adjacent event bins.
To achieve continuous predictions of high-frame-rate optical flow, it is necessary to obtain high-frame-rate event representations. Every input requires a uniform format for event representation. However, the Voxel Grid [17] represents the first and last channels differently as compared to the others, as shown in Fig. 3. The event interpolation time range utilized for the first and last channels in Voxel Grid is half of that used for the other channels, leading to inconsistency, which is adverse to continuous predictions. As a solution, we propose Unified Voxel Grid. In contrast to Voxel Grid, we fix the time interval $\tau$ for each bin. And each bin is created through the interpolation of events from the adjacent $\tau$ time-step ($t_{b}-\tau<e_{i}[t]<t_{b}+\tau$). The definition of UVG is as follows:
$Bin_{b}(x,y)=\sum_{i}p_{i}k_{b}\left(x-x_{i}\right)k_{b}\left(y-y_{i}\right)k_{b}\left(\frac{t-t_{b}}{\tau}\right)$ (2)
$UVG=concat([Bin_{0},Bin_{1},...,Bin_{B-1}]).$
This design makes it possible to obtain the current representation of time for each $\tau$ ($7.14ms$ for $15$ channels with Unified Voxel Grid on DESC), while the Voxel Grid must wait for the entire period of $(B-1)\times\tau$ ($100ms$ on DSEC) to generate a complete event representation. This low-latency input can be coupled with our low-latency time-dense optical flow estimation architecture for tracking the optical flow at each $\tau$ timestep as it arrives which is essential for high-frame-rate optical flow estimation.
III-B Event Anytime Flow Estimation Framework
@@
Serialized, low-latency input and output. The UVG inputs are sequentially fed into the network, which contributes to the low latency at the data entry level. During the inference phase, after entering an event bin (excluding the first bin used for initialization), the current optical flow can be directly predicted using the encoder and the stacked SMR module. During the training phase, in order to facilitate end-to-end training, we follow the frame rate of the dataset and input a full period of UVG at once as a training sample. Following the UVG input of $N,B,H,W$, the UVGs of different bins (i.e. with different channels) are converted to the batch dimension and concatenated (resulting in a dimension of $N{\times}B,H,W$) before being passed through the encoder. Once the feature maps are obtained, the corresponding feature maps of different bins are sequentially input into the stacked SMR module, which outputs time-dense optical flow ($\mathbf{V}_{0,1}\to\mathbf{V}_{0,B-1}$).
Spatiotemporal Motion Recurrent (SMR) module.
Figure 4: The proposed Spatiotemporal Motion Recurrent (SMR) module. The red dashed box encloses the output of the SMR module in the preceding feature pyramid layer, and the purple box surrounds the output of the current feature pyramid layer’s SMR module, which will transfer to the SMR module of the next layer in the feature pyramid.
In order to predict time-dense optical flow and achieve high accuracy through iterative refinement, we propose the stacked SMR modules, whose structure is illustrated in Fig. 4. In the vertical direction, as illustrated in Fig. 2 and Fig. 4, SMR performs an iterative refinement of optical flow at the current moment in the spatial dimension. The input of this module includes the hidden state of the lower-resolution SMR module from the previous level, the optical flow output, and the feature inputs of the current resolution level. First, we warp the input feature $f_{j}^{i}$ using the rough optical flow estimated from the previous level to obtain the post-warp features $\hat{f_{j}^{i}}$. Next, the optical flow $\mathbf{V}_{0,j}^{i-1}$ from the previous level, the warped features, and the upsampled hidden state $Ht_{j}^{i-1}$ from the previous level are concatenated, and used as input for the Convolutional Gated Recurrent Unit (ConvGRU) [45]. Subsequently, the hidden state output of the current ConvGRU is used by the Flow Head to predict the residual optical flow. We employ two convolutional layers for the Flow Head, which is the same as RAFT [25]. The refined optical flow $\mathbf{V}_{0,j}^{i}$ is achieved by adding the predicted residual optical flow $\Delta\mathbf{V}_{0,j}^{i}$ to the previously predicted optical flow $\mathbf{V}_{0,j}^{i-1}$. The current level’s predicted optical flow and hidden state are also inputted to the next level in the same way, for further refinement. This continues until the final level predicts the ultimate refined optical flow.
In the horizontal direction in Fig. 2, SMR continuously estimates optical flow at each time step. In general, SMR serves as a motion-updating module in both spatial and temporal dimensions. Within the same layer in the horizontal direction, SMR modules leverage shared weights, while in the vertical direction, different SMR weights are used due to varying input resolutions and channel numbers. In general, the number of SMRs is equal to the number of levels for spatial refinement of the optical flow estimation.
@@
Time-dense feature warping. Unlike other methods [39, 38] that rely on the assumption of uniform motion between two ground truth flow, our approach can handle variations in motion speeds throughout the entire UVG period. In each warping step, we use the dense temporal optical flow estimation of the previous level to warp the features of each time step. This structure guarantees consistent prediction patterns of optical flow for each resolution level and time step (SMR modules at the same level are using shared weights), implicitly providing SMR modules with the ability to predict optical flow at the current time step. By implementing this approach, we only need to supervise the final output $\mathbf{V}_{0,B-1}$ for implicit monitoring of optical flow at intermediate time steps.
III-C Supervision
@@
Following E-RAFT [11], we also used L1 loss as the loss function. Specifically, we calculate the L1 distance between the ground-truth optical flow and the last optical flow prediction $\mathbf{V}_{0,B-1}$ of our model as the final loss function. The loss function is defined as follows:
@@
Due to network structure design (see explanation in Sec. III-B in Time-dense feature warping), we don’t need to supervise the optical flow prediction at intermediate time steps. In this way, we overcome the frame rate limitation of the dataset and achieve low-latency, high-frame-rate, and high-precision event optical flow estimation.
Figure 5: RFWL vs. FWL. The first row shows the event count images generated by the original event stream of DSEC test sequences [12], while the second row shows the event count images generated by the motion-compensated event stream using our flow estimation, which we name MC frames. Each image is annotated with the variance and sum of pixel values (corresponding to the number of events detected in the respective image regions). The last row presents the FWL and RFWL of the motion-compensated frames (MC frames) for distinct columns.
III-D Rectified Flow Warp Loss
@@
Motion compensation [31] of events involves aligning all events to a reference time using estimated event optical flow. Specifically, for every event, using the flow value of the event position and the time difference between the event occurrence time and the reference time, we can calculate the position of the event point at the reference time. This process is referred to as motion compensation. The motion-compensated event count image is called Motion-Compensated (MC) event frame.
@@
Given events $E=(\left\{\left(x_{i},y_{i},t_{i},p_{i}\right)\right\}_{i\in[1,N]})$, per-pixel flow estimation $\mathbf{V}$ and a reference time $t^{\prime}$, the MC frame is defined as follow:
$I(E,\phi)=\left(\begin{array}[]{l}x_{i}^{\prime}\\ y_{i}^{\prime}\end{array}\right)=\left(\begin{array}[]{l}x_{i}\\ y_{i}\end{array}\right)+\left(t^{\prime}-t_{i}\right)\mathbf{V}(x_{i},y_{i}).$ (4)
@@
Accurate flow estimation ensures that events generated by the same edge are aligned to the same pixel location, resulting in higher contrast in the MC frame than in the original event frame. Stoffregen et al. [46] utilized this principle to propose a Flow Warp Loss (FWL), which can assess the accuracy of event optical flow estimation in an unsupervised way.
$\mathrm{FWL}:=\frac{\sigma^{2}(I(E,\mathbf{V}))}{\sigma^{2}(I(E,0))}$ (5)
@@
In theory, a more accurate estimation of optical flow results in a clearer, higher contrast, and larger variance for the image after motion compensation. This is because events occurring at the same location in the scene are aligned to the same point. Therefore, a more accurate prediction of optical flow leads to a larger FWL. In general, the image tends to become sharper after motion compensation, indicating an FWL value greater than $1$. However, in practice, we observe that FWL can be smaller than $1$, yet the MC frame is clearly sharper than the original one, as shown in Fig. 5. We have identified the same phenomena (FWL <1) in another work [32]. After analysis, we discover that motion compensation has warped some events out of the image, resulting in fewer events in the MC frame compared to the original frame. Thus, we propose a Rectified Flow warp Loss (RFWL), which normalizes image brightness based on the total event count (i.e., the sum of pixel values) in the event frames before and after motion compensation. The formulation is given as follows:
$\mathrm{RFWL}:=\sigma^{2}(\frac{I(E,\mathbf{V})}{\sum I(E,\mathbf{V})})/\sigma^{2}(\frac{I(E,0)}{\sum I(E,0)}).$ (6)
As shown in Fig. 5, FWL is inadequate at precisely assessing the accuracy of optical flow estimation. MC frames, as shown in columns 1, 2, and 4 exhibit sharper resolution compared to the original event frames, whereas the FWL values remain less than $1$. Interestingly, the MC frame in column 3, which has a significantly sharper resolution than the original event frame, yields an FWL value of only $1.12$. The FWL values of these examples are inconsistent with the actual observational results. Conversely, our proposed RFWL can precisely indicate the relative sharpness of MC frames in relation to the original event frames.
IV Experiments
@@
Our model underwent evaluation on the DSEC [12], MVSEC [14], and self-collected EVA-FlowSet. Sec IV-A offers an introduction to the aforementioned datasets. Sec. IV-B presents the implementation details of our approach. Sec. IV-C demonstrates the evaluation of our model using the regular flow evaluation prototype. Sec. IV-D reveals the evaluation results of our model for time-dense optical flow. Sec. IV-E encompasses ablation studies.
IV-A Datasets

### v3

@@ 3 Method
In this paper, we propose an Event Anytime Flow Estimation (EVA-Flow) framework, as shown in Fig. 2, which achieves high accuracy, time-dense flow estimation, and low latency input and output with supervision solely dependent on low frame rate optical flow. We first put forward a Unified Voxel Grid (UVG) representation in Sec. 3.1, which generates event representations with low latency. These UVG bins are sequentially fed into EVA-Flow, where high-frame-rate optical flow estimations are generated with low latency using a Spatiotemporal Motion Recurrent (SMR) module, which is detailed in Sec. 3.2. Then, the details of the loss function and the supervision regime are described in Sec. 3.3. Finally, we propose a new criterion to evaluate the reliability of intermediate high-frame-rate optical flow in Sec. 3.4.
{adjustwidth}
[…]
Figure 2: Architecture of our Event Anytime Flow Estimation (EVA-Flow) framework (a) and SMR module (b). $f^{i}$ denotes the features of the $i$-th level in a feature pyramid, while $f^{i}_{j}$ denotes the $j$-th event bin’s feature in $j^{i}$. $F_{0,i}$ denotes flow prediction from time $t_{0}$ to $t_{i}$. The red dashed box indicates the previous layer’s SMR module output, while the purple box shows the current layer’s output, destined for the next layer’s SMR module.
3.1 Event Representation: Unified Voxel Grid
@@
The event representation is crucial for the event-based model Baldwin et al. (2023); Gehrig et al. (2019). Voxel Grid Zhu et al. (2019) is a commonly used event representation that utilizes the temporal dimension of event data, and its effectiveness for optical flow estimation tasks has been verified in many studies Zhu et al. (2019); Gehrig et al. (2021); Wu et al. (2024); Gehrig et al. (2024). The principle of the Voxel Grid is to discretize the spatially sparse and temporally-continuous information of events by averaging over time, and then utilize bilinear interpolation in both spatial and temporal dimensions to obtain a tensor representation of the event data in a voxel form. The tensor has a dimension of $B{\times}H{\times}W$, where $B$ represents the number of time steps, and its size determines the sampling accuracy of the event data in the temporal dimension. Each channel of the Voxel Grid can be viewed as an event representation at a specific time. Given a set of events $\left\{\left(x_{i},y_{i},t_{i},p_{i}\right)\right\}_{i\in[1,N]}$ and $B$ bins to discretize the time dimension, the definition of Voxel Grid is depicted as follows:
@@
Figure 3: The proposed Unified Voxel Grid compared with Voxel Grid Zhu et al. (2019). Different colored lines demonstrate the time range of events used for interpolation in different bins, as well as the interpolation weights corresponding to those events.
Continuous prediction of high-frame-rate optical flow necessitates corresponding high-frequency event representations. A critical limitation of the standard Vooxel Grid, illustrated in Fig. 3, is that generating the complete grid requires waiting for all events within the entire time interval $T$, making it inherently batch-based and unsuitable for low-latency applications. Furthermore, the effective temporal support for the first ($b=0$) and last ($b=B-1$) bins is often halved, leading to representational inconsistency.
To address these limitations, particularly the latency issue, we propose the Unified Voxel Grid (UVG). UVG enables sequential, low-latency event representation. Instead of waiting for the full duration $T$, UVG allows for the computation of an event representation centered around a specific time $t_{current}$ using only events within a localized temporal window. We achieve this by defining each bin $Bin_{b}$ consistently with a fixed temporal support $2\tau$, centered at its timestamp $t_{b}$. The representation for a bin is calculated using events $e_{i}$ falling within $t_{b}-\tau<t_{i}<t_{b}+\tau$:
$Bin_{b}^{UVG}(x,y)=\sum_{i}p_{i}k_{b}\left(x-x_{i}\right)k_{b}\left(y-y_{i}\right)k_{b}\left(\frac{t-t_{b}}{\tau}\right)$ (2)
This design makes it possible to obtain the current representation of time for each $\tau$ ($5\,ms$ for $21$ channels with Unified Voxel Grid on DESC), while the Voxel Grid must wait for the entire period of $T$ ($100\,ms$ on DSEC) to generate a complete event representation. This low-latency input can be coupled with our low-latency time-dense optical flow estimation architecture for tracking the optical flow at each $\tau$ timestep as it arrives which is essential for high-frame-rate optical flow estimation.
3.2 Event Anytime Flow Estimation Framework
@@
Serialized, low-latency input and output. The UVG inputs are sequentially fed into the network, which contributes to the low latency at the data entry level. During the inference phase, after entering an event bin (excluding the first bin used for initialization), the current optical flow can be directly predicted using the encoder and the stacked SMR module. During the training phase, to facilitate end-to-end training, we follow the frame rate of the dataset and input a full period of UVG at once as a training sample. Following the UVG input of $N,B,H,W$, the UVGs of different bins (i.e. with different channels) are converted to the batch dimension and concatenated (resulting in a dimension of $N{\times}B,H,W$) before being passed through the encoder. Once the feature maps are obtained, the corresponding feature maps of different bins are sequentially input into the stacked SMR module, which outputs time-dense optical flow ($\mathbf{F}_{0,1}\to\mathbf{F}_{0,B-1}$).
Spatiotemporal Motion Recurrent (SMR) module. To achieve accurate and time-dense optical flow estimation, we propose the Spatiotemporal Motion Recurrent (SMR) module, illustrated in Fig. 2. The core idea of SMR is to employ a hierarchically stacked structure for coarse-to-fine spatial refinement combined with ConvGRU-based temporal recurrence. This design inherently promotes stable convergence and high accuracy. Specifically, the coarse-to-fine spatial processing provides robust initial estimates and guides the refinement process across scales, simplifying the overall estimation task. Concurrently, the iterative residual updates within each level encourage gradual convergence by successively correcting the flow estimate. Moreover, the ConvGRU architecture Ballas et al. (2015), with its gating mechanisms, ensures stable learning and propagation of temporal motion information, effectively capturing motion dynamics over time. The process within each SMR unit is governed by:
$\hat{f_{j}^{i}}=\textbf{Warp}(f_{j}^{i},\mathbf{F}_{0,j}^{i-1})$ (3)
$Ht_{j}^{i}=\textbf{ConvGRU}(Ht_{j-1}^{i},concat(\hat{f_{j}^{i}},\text{Up}(\mathbf{F}_{0,j}^{i-1}),\text{Up}(Ht_{j}^{i-1}))$
$\mathbf{F}_{0,j}^{i}=\mathbf{F}_{0,j}^{i-1}+\textbf{FlowHead}(Ht_{j}^{i})$
where superscript $i$ denotes the feature pyramid level (from coarse to fine) and subscript $j$ indicates the temporal processing step.
Structurally, the SMR module operates along two primary dimensions:
Vertical Dimension (Spatial Coarse-to-Fine Refinement): At any given time step $j$, the SMR modules are stacked vertically across different spatial resolutions (pyramid levels $i$). Information propagates from coarser levels ($i-1$) to finer levels ($i$). Specifically, the input features $f_{j}^{i}$ at level $i$ are warped using the upsampled flow estimate $\mathbf{F}_{0,j}^{i-1}$ from the coarser level. This aligned feature $\hat{f_{j}^{i}}$, along with the upsampled flow and the upsampled hidden state $Ht_{j}^{i-1}$, are fed into the ConvGRU unit. The resulting hidden state $Ht_{j}^{i}$ is then processed by a Flow Head (implemented with two convolutional layers) to predict a residual flow, which updates the coarser flow estimate $\mathbf{F}_{0,j}^{i-1}$ to yield the refined flow $\mathbf{F}_{0,j}^{i}$ at the current level. This hierarchical refinement progressively enhances spatial accuracy. Importantly, the optical flow $\mathbf{F}_{0,j}^{i}$ for the current time step $j$ is predicted using only the features $f_{j}^{i}$ and state information available up to that time step. This enables low-latency flow estimation, as the output for time $j$ can be generated as soon as the corresponding event features are processed.
Horizontal Dimension (Temporal Recurrence): Within each spatial level $i$, the SMR module processes information sequentially across time steps $j$. The ConvGRU Ballas et al. (2015) plays a crucial role here, utilizing its internal hidden state ($Ht_{j-1}^{i}$ from the previous time step) to update the current hidden state $Ht_{j}^{i}$. This recurrence mechanism maintains temporal consistency and allows the module to model motion dynamics over time. SMR modules within the same spatial level $i$ (horizontally) share weights across different time steps $j$, promoting efficient learning of temporal patterns. Conversely, SMR modules at different spatial levels $i$ (vertically) use distinct weights to accommodate varying feature resolutions and channel dimensions.
@@
Time-dense feature warping. Unlike other methods Wu et al. (2024); Liu et al. (2023) that rely on the assumption of linear motion between two ground-truth flows, our approach can handle non-linear motion throughout the entire UVG period. As depicted in Figure 2, during each warping iteration, the dense temporal optical flow from the preceding level is employed to warp the features of the current level at each time step. This structure guarantees consistent prediction patterns of optical flow for each resolution level and time step (SMR modules at the same level are using shared weights), implicitly providing SMR modules with the ability to predict optical flow at the current time step. By implementing this approach, we only need to supervise the final output $\mathbf{F}_{0,B-1}$ for implicit monitoring of optical flow at intermediate time steps.
3.3 Supervision
@@
Following E-RAFT Gehrig et al. (2021), we also used L1 loss as the loss function. Specifically, we calculate the L1 distance between the ground-truth optical flow and the last optical flow prediction $\mathbf{F}_{0,B-1}$ of our model as the final loss function. The loss function is defined as follows:
@@
Due to network structure design (see explanation in Sec. 3.2 in Time-dense feature warping), we don’t need to supervise the optical flow prediction at intermediate time steps.
{adjustwidth}
[…]
Figure 4: RFWL vs. FWL. The top row displays event count images from the original DSEC-Flow dataset Gehrig et al. (2021), with the second row presenting the corresponding Images of Warped Events (IWE). Variance and sum are indicated for each image. The bottom row shows the FWL and RFWL for the IWEs.
3.4 Rectified Flow Warp Loss
@@
Motion compensation, as described in Gallego et al. (2018), entails aligning all events to a reference time by utilizing the estimated event optical flow. Specifically, for each event, the flow value at the event’s position and the time difference from the event’s occurrence to the reference time are used to determine the event’s position at the reference time (see Equation. 5). Henceforth, we refer to the motion-compensated event count image as the Image of Warped Events (IWE).
@@
Given events $E=(\left\{\left(x_{i},y_{i},t_{i},p_{i}\right)\right\}_{i\in[1,N]})$, per-pixel flow estimation $\mathbf{F}$ and a reference time $t^{\prime}$, the IWE is defined as follow:
$I(E,\mathbf{F})=\left(\begin{array}[]{l}x_{i}^{\prime}\\ y_{i}^{\prime}\end{array}\right)=\left(\begin{array}[]{l}x_{i}\\ y_{i}\end{array}\right)+\left(t^{\prime}-t_{i}\right)\mathbf{F}(x_{i},y_{i}).$ (5)
@@
Accurate flow estimation ensures that events generated by the same edge are aligned to the same pixel location, resulting in higher contrast in the IWE than in the original event frame. Stoffregen et al. Stoffregen et al. (2020) utilized this principle to propose a Flow Warp Loss (FWL), which can assess the accuracy of event optical flow estimation in an unsupervised way.
$\mathrm{FWL}:=\frac{\sigma^{2}(I(E,\mathbf{F}))}{\sigma^{2}(I(E,0))}$ (6)
@@
In theory, a more accurate estimation of optical flow results in a clearer, higher contrast, and larger variance for the image after motion compensation. This is because events occurring at the same location in the scene are aligned to the same point. Therefore, a more accurate prediction of optical flow leads to a larger FWL. In general, the image tends to become sharper after motion compensation, indicating an FWL value greater than $1$. However, in practice, we find that the FWL can be less than 1, yet the IWE remains visibly sharper than the original, as depicted in Figure 4. This phenomenon, where FWL of IWE is less than 1, has also been observed in the study by Shiba et al. Shiba et al. (2022). Upon analysis, we determined that motion compensation has caused some events to be warped outside the image boundaries, leading to a reduced event count in the IWE compared to the original event frame. Thus, we propose a Rectified Flow warp Loss (RFWL), which normalizes image brightness based on the total event count (i.e., the sum of pixel values) in the event frames before and after motion compensation. The formulation is given as follows:
$\mathrm{RFWL}:=\sigma^{2}(\frac{I(E,\mathbf{F})}{\sum I(E,\mathbf{F})})/\sigma^{2}(\frac{I(E,0)}{\sum I(E,0)}).$ (7)
As depicted in Fig. 4, the FWL fails to accurately assess the precision of optical flow estimation. The IWEs in columns 1, 2, and 4 display higher sharpness than the original event frames, yet the corresponding FWL values are less than $1$. Notably, the IWE in column 3, which exhibits a markedly sharper resolution than the original, has an FWL value of $1.12$. In contrast, our proposed Rectified Flow Warp Loss can accurately reflect the relative sharpness of IWEs compared to the original event frames.
4 Experimental Results
@@
Our model underwent evaluation on the DSEC Gehrig et al. (2021), MVSEC Zhu et al. (2018), and self-collected EVA-FlowSet. Sec 4.1 offers an introduction to the aforementioned datasets. Sec. 4.2 presents the implementation details of our approach. Sec. 4.3 demonstrates the evaluation of our model using the regular flow evaluation prototype. Sec. 4.4 reveals the evaluation results of our model for time-dense optical flow. Sec. 4.5 encompasses ablation studies.
4.1 Datasets